A ball is dropped from rest and feels air resistance as it falls. Which of the graphs in Fig. Q5.26 best represents its vertical

velocity component as a function of time?



1
Expert's answer
2021-11-29T11:41:22-0500


air resistance depends on v2, so equation of motion:

"ma=mg-kv^2"

"mv'=mg-kv^2"


"\frac{mdv}{mg-kv^2}=dt"


solution is:

"arctan(k_1v)=k_2t+k"

or

"v=tan(k_2t+k)\/k_1"


Answer: c